OpenStudy (anonymous):

why isn't 21^3*33*25^14 is a prime factorization?

jimthompson5910 (jim_thompson5910):

because 21 and 25 aren't prime numbers

jimthompson5910 (jim_thompson5910):

oh and 33 as well

OpenStudy (anonymous):

how about 105, is it a prime???

jimthompson5910 (jim_thompson5910):

no since it's divisible by 3

jimthompson5910 (jim_thompson5910):

105/3 = 35

OpenStudy (anonymous):

but 145 is a prime?

jimthompson5910 (jim_thompson5910):

no, it's divisible by 5

jimthompson5910 (jim_thompson5910):

since it ends with a 5

OpenStudy (anonymous):

all right, got it. thanks for helping...

OpenStudy (unklerhaukus):

\[21^3*33*25^{14}\] \[=3^{3}*7^{3}*3*11*5^{28}\] \[=3^{4}*5^{28}*7^{3}*11\] now in prime factorization

OpenStudy (anonymous):

145 is not prime